I have two problems that I just can't seem to get worked out. I submitted one before but the wrong numbers were used in the equation by the solver. These problems are due tonight by midnight. Any help give will be gravely appreciated.
Question 1
5x - 10x^2 < 0
Question 2
x^2 - x > 0
Solve the inequality, state the solution set in inteval notation and graph it.

Answer by stanbon(75887)

5x-10X^2<0
1st solve the equality 5x-10x^2=0
5x(1-2x=0
x=0 or x=1/2
Draw a number line and plot x=0 and x=1/2 on it
You now have three sections on the line; they
are x<0; 01/2
Check a value in each section in the original
inequality to see if that value makes the equality
true or false.
For example try x=-10 to check the section x<0
Substituting you get:
5(-10)-10(-10)^2<0
-50-1000<0
That's true so all x<0 are part of the solution.
Check also x=1/4, and x=10 to see if either of
the other two sections contain solutions.
Cheers,
Stan H.